In the centre of a rectangular lawn of dimensions 50 m × 40 m, a rectangular
pond has to be constructed so that the area of the grass surrounding the pond
would be 1184 m2 [see Fig. 4.1]. Find the length and breadth of the pond.

Grade:10

2 Answers

Ashwin
9 years ago
Length of park = L = 50
Breadth of park = B= 40 
area of rectangular field =L x B= 50 x 40 = 2000m2
 therefore , area of rect pond = area of park – area of grass2000 - 1184 = 816m2
=> l x b = 816 m2
and length and breadth can have many values but if you need exact value you need to do more calculations ….. i am in a hurry i cant do them will post soon
Sourav
6 years ago
50*40-(50-2x) (40-2x)=11842000-(50-2x)(40-2x)=1184100x+80x-4x2=11844x2-180x+1184=04(x2-45x+296)=0(x-37) (x-8)=0X=37, 8Since x =37 is no possible , as otherwise the length of pond will be negative.X=8 is the required solution Length= 50-2*8=34 mAnd breadth =40-2*8=24m